A table of values for the equation y=f(x) contains an x value of -6 and a corresponding y value of 12, as does a table of values for the equatio y=g(x). which of these must be the solution to the equation f(x)=g(x)
Answer by Edwin McCravy(20056)


Let's take an example where that is the case:

y = f(x) = -3x-6 has this table of values:

 x | y
-3 | 3
 1 |-9
-6 |12
-5 | 9 

y = g(x) = 4x+36 has this table of values

 x | y
-8 | 4
-6 |12
 1 |40
-9 | 0

As we see above, the table of values for the equation y = f(x) 
contains an x value of -6 and a corresponding y value of 12, 
as does the table of values for the equation y = g(x). 

What we want to know is which of these must be the 
solution to this equation: 

 f(x) = g(x)

Substitute what f(x) equals for f(x) and what g(x) equals
for g(x) in that equation, like this 

-3x-6 = 4x+36

Find the solution to that, then answer the question asked.

Edwin